OpenStudy (anonymous):

solve: what are the x-intercepts of the graph of the equation y − x2 = −4x − 5?

OpenStudy (btaylor):

put it in simplified form: y = x^2-4x-5 factor into 0 = (x-5)(x+1) set both factors equal to zero, solve for x x = 5 and -1, so the x-intercepts are (5,0) snd (-1,0)

OpenStudy (anonymous):

^

OpenStudy (anonymous):

oh okay! thanks! im not very bright in algebra...some things i get and some i dont..thanks for your help! :)